OpenStudy (anonymous):

Which point lies inside the circle with equation (x - 2)2 + (y + 3)2 = 4? 0,-3 2,-4 2,0 2,-5

OpenStudy (danjs):

center (h,k) = (2,-3)

OpenStudy (danjs):

radius = 2

OpenStudy (danjs):



OpenStudy (danjs):

2 of the choices lie ON the circle, 1 of the choices is IN the circle 1 of the choices is OUTSIDE the circle

OpenStudy (anonymous):

but there is no (2,-3) listed for the answers

OpenStudy (danjs):

that is the center of the circle

OpenStudy (danjs):

you want one that is INSIDE the circle

Parth (parthkohli):

Addition: If a point \((x_0 ,y_0 ) \) lies inside a circle, then its distance from the centre is less than the radius.\[\sqrt{(x_0 - h)^2 + (y_0 - k)^2} < r\]\[(x_0 - h)^2 + (y _0 - k)^2 < r^2\]

OpenStudy (danjs):

if you say (x-2)^2 + (y+3)^2 < 4

OpenStudy (freckles):

If you have for a point (x, y) that (x-2)^2 +(y+3)^2 <4, then the point (x, y) is inside. If you have for a point (x, y) that (x-2)^2 +(y+3)^2 >4 then (x, y) is outside the circle. If you have for a point (x, y) that (x-2)^2 +(y+3)^2 =4 then the point (x, y) is on the circle.

Parth (parthkohli):

In other words, all the points satisfying \((x - h)^2 + (y - k)^2 < r^2\) will lie inside the circle.

OpenStudy (alekos):

You would be better off doing a sketch on graph paper and solving geometrically.

OpenStudy (danjs):

right, i did that above there

OpenStudy (anonymous):

(2 , -3)